OpenStudy (anonymous):

What is the equation of the line that passes through the points (-2, 1) and (1, 10)?

OpenStudy (anonymous):

I'm starting to feel this line pass two points question comes every minute here :)

OpenStudy (anonymous):

the equation is y-y1/x-x1=m. where m is the slope .. and m =y2-y1/x2-x1

OpenStudy (anonymous):

haha well i just don't get which formula to use

hero (hero):

Insert each point into the formula y = mx + b: 1 = -2m + b 10 = m + b Isolate b in each equation: b = 1 + 2m b = 10 - m Set b = b 1 + 2m = 10 - m solve for m 2m + m = 10 - 1 3m = 9 m = 9/3 m = 3 Find b: b = 1 + 2m b = 1 + 2(3) b = 1 + 6 b = 7 Therefore y = 3x + 7

hero (hero):

You only need to remember one formula for this: y = mx + b

OpenStudy (anonymous):

okay .. m=10-1/1+3..

OpenStudy (anonymous):